OpenStudy (anonymous):

can anyone help me solve this PLEASE~!! The volume of a gas varies directly as its temperature and inversely as its pressure. AT a temperature of 100 Kelvin and a pressure of 15 kilograms per square meter, the gas occupies a volume of 20 cubic meters. Find the volume at a temperature of 150 Kelvin and a pressure of 30 kilograms per square meter.

OpenStudy (anonymous):

anyone ???

OpenStudy (ganpat):

V = (K) * T/P K is some constant.. can u go further ??

OpenStudy (anonymous):

i have that but i dont know which numbers to plug in for what

OpenStudy (ganpat):

ohkey, Now consider 1st.. 20 = K (100/ 15) K = 3.. SO, For second values... V = K (T/P) V = 3 (150/30) = 15 cubic meter...

OpenStudy (anonymous):

oh ok that makes much more sense